# Effects of dynamical friction on perturbations for evolving dark energy

### v1

@@ II.1 A comment on dynamical friction
where $t$ is the cosmic time, $c$ the speed of light, $G$ Newton’s constant, $\Lambda$ the value of the cosmological constant and $M_{\rm m}$ the mass of the dark matter perturbation. This expression has been explicitly written down in Del Popolo et al. (2019) to derive a mass-temperature relation for clusters in $\Lambda$CDM and modified gravity models.
Following Pace et al. (2019), we can easily extend the previous expression to an arbitrary dark energy model described by an equation-of-state parameter $w_{\rm de}$ and express the angular momentum contribution in terms of the shear $\sigma$ and rotation $\omega$ term
@@
$M_{\rm m}=\frac{4\pi}{3}\bar{\rho}_{\rm m}(1+\delta)R^{3}\,,\quad M_{\rm bck,de}=\frac{4\pi}{3}\bar{\rho}_{\rm de}R^{3}\,.$ (9)
In the previous expressions, $\bar{\rho}_{\rm m}$ and $\bar{\rho}_{\rm de}$ are the background matter and dark energy density, respectively and $M_{\rm bck,de}$ represents the amount of dark energy included in a sphere of radius $R$. This expression can be further generalised to take into account dark energy perturbations as in Pace and Schimd (2022)
$\frac{\mathrm{d}^{2}R}{\mathrm{d}t^{2}}=$ $\,-\eta\frac{\mathrm{d}R}{\mathrm{d}t}-\frac{GM_{\mathrm{m}}}{R^{2}}-\frac{GM_{\mathrm{de}}}{R^{2}}(1+3s_{\rm eff})$
$\,-\frac{GM_{\rm bck,de}}{R^{2}}(1+3w_{\mathrm{de}})-\frac{\sigma^{2}-\omega^{2}}{3}R\,,$ (10)
where
$M_{\mathrm{de}}=\frac{4\pi}{3}\bar{\rho}_{\mathrm{de}}(a)\delta_{\mathrm{de}}R^{3}\,,$ (11)
is the mass associated to dark energy perturbation. Note that this equation does not say anything whether also dark energy experiences dynamical friction.
As shown in Appendix B, Eq. (8) leads to
@@
which contains the two additional terms $-\tfrac{3\eta}{H}(1+\delta)$ and $\frac{\eta}{H}\delta^{\prime}$. The latter is a proper friction term. At the linear level, it reads $-\tfrac{3\eta}{H}$ and leads to conceptual problems. It acts as a source term such that the particular solution is, for a model where $\eta=\eta_{0}H$ and for an EdS model, $\delta=-2\eta_{0}$. This implies that for $a\to 0$ an overdensity originates from an underdensity, or said in other terms, primordial perturbations are small, but do not go to zero as expected.
As noted also in Carrilho et al. (2022), this is due to the fact that dynamical friction only depends on the peculiar velocity of dark matter and that while for the overdensity we use comoving coordinates, for the evolution of the radius we use physical coordinates. In particular the radius of the perturbation is the physical radius. When we go from the evolution of the density perturbation $\delta$ to that of $R$, using the same procedure outlined in Appendix B, the equation for the latter is (see also Carrilho et al., 2022)
@@
This equation contains two terms originating from friction: one contribution dampens the collapse ($-\eta\,\mathrm{d}R/\mathrm{d}t$) and the other ($\eta HR$) enhances it. It is this term, $\eta HR$, which exactly cancels the term $3\eta(1+\delta)/H$ in Eq. (II.1) and thus curing the aforementioned problem, i.e. that perturbations are sourced, rather than being slowed down by friction.
This can be also seen by defining the comoving variable $y\equiv\tfrac{R}{R_{i}}\tfrac{a_{i}}{a}=\lambda\tfrac{R}{a}$, so that Eq. (II.1) becomes
$\ddot{y}+(2H+\eta)\dot{y}+\frac{1}{2}H^{2}\Omega_{\rm m}(a)\delta y+\frac{\sigma^{2}-\omega^{2}}{3}y=0\,,$ (14)
with $\delta$ the density contrast of the perturbation defined as
$\delta=[1+\delta_{i}]\left(\frac{a_{i}}{a}y\right)^{-3}-1\,.$ (15)
If we instead consider the original formulation of dynamical friction via the equation of motion of the physical radius $R$, Eq. (8), we arrive to
$\ddot{y}+(2H+\eta)\dot{y}+\frac{1}{2}H^{2}\Omega_{\rm m}(a)\delta y+\eta Hy+\frac{\sigma^{2}-\omega^{2}}{3}y=0\,,$ (16)
where, once again, we see a term arising from the mixture of comoving and physical quantities. With respect to Carrilho et al. (2022) we also took into account shear and rotation terms. From now on, we will neglect their contribution in our analysis and set them to zero, effectively considering an exact spherical symmetry, as we want to focus on the effects of friction and avoid an interplay among them and friction. The effects of shear and rotation on structure formation have been studied before for both smooth and clustering dark energy models and for different implementations, as already discussed in Sect. II. We will not delve into them again, but we refer the reader to Del Popolo et al. (2013a, b, c); Pace et al. (2014b); Reischke et al. (2016); Pace et al. (2017a); Reischke et al. (2018); Pace et al. (2019); Pace and Schimd (2022) for details.
In Sect. III we detail how to calculate the time dependence of $\eta$ with respect to time. Here we simply summarise our findings. For the dynamical friction model of Antonuccio-Delogu and Colafrancesco (1994) we obtain $\eta=\eta_{0}\,H$ with $\eta_{0}$ constant and for dissipative pressure friction we find $\eta=\eta_{0}(\eta_{1}+\eta_{2}|P/P_{0}|^{\Gamma})\,H_{0}$, with $\eta_{1}$, $\eta_{2}$ and $\Gamma$ all constant.
III Theoretical models of friction
@@
III.1 Evolution of dynamical friction
Here, we follow closely the works by Antonuccio-Delogu and Colafrancesco (1994) and Del Popolo (2009), to which we refer for more details.
In a scenario with hierarchical structure formation, structures of size $R$ form around the local maxima of the primordial density field smoothed over a scale of size $R$ Bardeen et al. (1986); Bond (1988); Colafrancesco et al. (1989). Assuming an initial Gaussian density field, the comoving averaged number density of peaks with central height $\nu\equiv\delta(z)/\sigma(M,z)$, with $\delta$ the overdensity and $\sigma$ the variance of the field, is given by
$\mathcal{N}_{\rm pk}(\nu|R)\,\mathbf{d}\nu=\frac{1}{(2\pi)^{2}R_{\ast}^{3}}G(\nu,\gamma\nu)\,e^{-\nu^{2}/2}\,,$ (17)
where the spectral parameters are $R_{\ast}=\sqrt{3}\sigma_{1}/\sigma_{2}$ and $\gamma=\sigma_{1}^{2}/(\sigma_{0}\sigma_{2})$, the function $G$ is given in Eq. (4.4) of Bardeen et al. (1986) and $\sigma_{l}$, with $l\in\{0,1,2\}$, are the spectral moments defined as
$\sigma_{l}^{2}(R,z)=\frac{1}{2\pi^{2}}\int\,\mathrm{d}k\,k^{2(1+l)}P(k,z)W(kR)\,,$ (18)
and $W(kR)$ is a filtering function.
In this system, a test particle experiences a gravitational field which consists of a part associated to the smooth global mass distribution and a stochastic one which originates from the particle number fluctuations and giving rise to a frictional force $-\eta\mathbf{v}$ where $\eta$ is the dynamical friction coefficient and $\mathbf{v}$ the macroscopic velocity.
Following Antonuccio-Delogu and Colafrancesco (1994), one can show that the peaks of the local density field with central height $\nu\geq\nu_{\rm c}$, with $\nu_{\rm c}$ a critical threshold, contribute to the dynamical friction. The number of these objects can be calculated under the condition that the peak radius $r_{\rm pk}(\nu\geq\nu_{\rm c})$ is negligible with respect to the average peak separation $n_{\rm a}(\nu\geq\nu_{\rm c})^{-1/3}$.
The values for $\nu_{\rm c}$ and the spectral moments will depend on the particular cosmological model considered, and the precise value of the spectral moments depend on the filtering scale $R$. Here we do not pretend to derive exact values for the quantities above for every individual model we investigate in this work. Our aim is merely to estimate the order of magnitude of the relevant quantities, in particular the dynamical friction coefficient $\eta$ and understand how this parameter impacts the evolution of linear and nonlinear structures in $\Lambda$CDM and dark energy models.
For a standard cosmological model with flat CDM spectrum Bond and Efstathiou (1984), Antonuccio-Delogu and Colafrancesco (1994); Del Popolo (2009) found $\nu_{\rm c}\approx 1.24$ and for a peak of mass $\approx 10^{9}\,h^{-1}\,M_{\odot}$ one obtains $\gamma\approx 0.442$ and $R_{\ast}\approx 45.2\,h^{-1}\,\text{kpc}$.
@@ III.1 Evolution of dynamical friction
where we used the fact that $m_{\rm a}n_{\rm a}=\rho_{\rm m}=\bar{\rho}_{\rm m}\Delta$, with $\Delta$ the average overdensity of the perturbation. In principle, $\Delta$ would depend on the virialization recipe and on the cosmological model. However, here for simplicity, we will assume it to be constant, of the order of $\mathcal{O}(100)$.
This result shows that $\eta\propto H$ and that the value of $\eta_{0}$ depends on the total number of peaks $N$ inside a protocluster, decreasing when $N$ increases. As discussed in Antonuccio-Delogu and Colafrancesco (1994), $N$ is considered to be approximately constant for clusters. For $N_{\rm tot}=10^{3}$ we find $\eta_{0}\approx 3\times 10^{-2}$, while for $N_{\rm tot}=10^{4}$ we have $\eta_{0}\approx 4\times 10^{-3}$. Regardless of the exact value of $\eta_{0}$, the main conclusion is that $\eta_{0}\ll 1$.
Note also that this result is not in contrast with previous works, such as Antonuccio-Delogu and Colafrancesco (1994), as there an EdS model was assumed. Indeed, for such a model, neglecting radiation, we have $H\propto a^{-3/2}$. In fact, Antonuccio-Delogu and Atrio-Barandela (1992) showed that it is possible to associate a relaxation timescale with dynamical friction in galaxy clusters and this is of the order of the Hubble time, i.e., $1/\eta\propto H^{-1}$.
Finally, rather than expressing $\eta$ in terms of the total number of peaks $N$, it is worth to use the mass of the perturbation $M$. Recalling that $M=m_{\rm a}N$, it is straightforward to obtain
@@
where we assume $m_{\rm a}=10^{9}\,h^{-1}\,M_{\odot}$. It is evident that $\eta_{0}$ is smaller for massive objects rather than low-mass perturbations.
Dynamical friction is a consequence of the fact that general relativity considers a pressure term that is absent in Newtonian dynamics. Dynamics is influenced by pressure and it is natural to follow this direction, i.e., friction originated by pressure to extend the standard dynamical friction model. We will describe this approach in the next section.
III.2 Dissipative pressure friction
Standard fluids can be grouped into two main classes: Newtonian and non-Newtonian. Naively, the latter is found when the friction coefficient, $\eta$, depends upon the variation of shear velocity with respect to the radial coordinate. Our approach is simplified by virtue of the cosmological principle, i.e., the velocity field does not depend upon radial coordinate. In this scenario, however, our friction coefficient is a function of the Hubble rate only. In other words, we do not emphasise a time-dependence of it, assuming the simplest stationary case. It is clearly possible to generalise this case, including the effects due to the dynamics of the universe. Following the standard treatment, one can assume that the friction coefficient is a function of temperature and/or of thermodynamic quantities such as pressure, entropy, enthalpy and so forth. In the simplest case, we can assume what follows
@@ III.2 Dissipative pressure friction
where $\eta_{1},\eta_{2}$ and $\Gamma$ are free real constants, with $P$ the total pressure of the universe and $P_{0}$ the total pressure today. The overall $H_{0}$ factor is just a normalization factor which makes the two coefficients $\eta_{1}$ and $\eta_{2}$ dimensionless. From the second Friedmann equation,
@@
At this stage, we distinguish two cases, as below reported.
@@
    reducing to our previous case, discussed above, with $\Gamma=\tfrac{1}{2}$, $\eta_{1}=0$ and $\eta_{2}=\eta_{0}$. This certifies that our previous approach turns out to be a suitable approximation that works well in a phase in which $H^{2}$ dominates over the dynamics of the universe, albeit it does not account for the whole stages of its evolution.
@@
Assuming that friction increases at large $z$, we require [^6] Remember that $\frac{\mathrm{d}}{\mathrm{d}x}|f(x)|=\frac{f(x)}{|f(x)|}\frac{\mathrm{d}f(x)}{\mathrm{d}x}$.
@@
recasting in terms of the jerk parameter, defined as $j\equiv-H^{-1}\dot{q}+2q^{2}+q$ Dunsby and Luongo (2016), with $q\equiv-1-\frac{\dot{H}}{H^{2}}$, the deceleration parameter.
Thus, since $\dot{H}>0$ and $j\geq 0$ Luongo et al. (2016); Capozziello et al. (2022), we have
@@
where $\eta_{1}$ is arbitrarily imposed to vanish as it appears unessential in determining $\eta$ and the values ±$\pm 1$ selected for normalizing $\eta_{2}$ and $\Gamma$.
Having the above, the case $\Gamma=0$ corresponds to a constant friction term, while the case $\eta_{2}=0$ implies a frictionless model. Consequently, we argue that
@@
    the second case described above implies a model where friction acts repulsively, i.e., the sign is compatible with the dark energy equation of state. We will not consider this case here as it does not represent the ordinary behaviour of friction on matter.
@@
IV Evolution of linear perturbations

### v2

@@ 2.1 A comment on dynamical friction
where $t$ is the cosmic time, $c$ the speed of light, $G$ Newton’s constant, $\Lambda$ the value of the cosmological constant and $M_{\rm m}$ the mass of the dark matter perturbation. This expression has been explicitly written down in [70] to derive a mass-temperature relation for clusters in $\Lambda$CDM and modified gravity models.
Following [52], we can easily extend the previous expression to an arbitrary dark energy model described by an equation-of-state parameter $w_{\rm de}$ and express the angular momentum contribution in terms of the shear $\sigma$ and rotation $\omega$ terms
@@
$M_{\rm m}=\frac{4\pi}{3}\bar{\rho}_{\rm m}(a)(1+\delta)R^{3}\,,\quad M_{\rm bck,de}=\frac{4\pi}{3}\bar{\rho}_{\rm de}(a)R^{3}\,.$ (9)
In the previous expressions, $\bar{\rho}_{\rm m}(a)$ and $\bar{\rho}_{\rm de}(a)$ are the background matter and dark energy density, respectively, and $M_{\rm bck,de}$ represents the amount of dark energy included in a sphere with radius $R$. [53] further extended the model to account also for dark energy perturbations.
It can be easily shown that Eq. (8) leads to
@@
which contains the two additional terms $-\tfrac{3\eta}{H}(1+\delta)$ and $\frac{\eta}{H}\delta^{\prime}$. The latter is a proper friction term. At the linear level, the first term reads $-\tfrac{3\eta}{H}$ and leads to conceptual problems. It acts as a source term so that the particular solution is, for a model where $\eta=\eta_{0}H$ and for an EdS model, $\delta=-2\eta_{0}$. This implies that for $a\to 0$ an overdensity originates from an underdensity, or to put it in other terms, primordial perturbations are small but do not go to zero as expected.
As also noted in [57], this is due to the fact that dynamical friction depends only on the peculiar velocity of dark matter and that while for the overdensity we use comoving coordinates, for the evolution of the radius we use physical coordinates. In particular, the radius of the perturbation is the physical radius. When we go from the evolution of the density perturbation $\delta$ to that of $R$, the equation for the latter is [see also 57]
@@
This equation contains two terms originating from friction: one contribution dampens the collapse ($-\eta\,\mathrm{d}R/\mathrm{d}t$) and the other ($\eta HR$) enhances it. It is the term $\eta HR$ that exactly cancels the term $3\eta(1+\delta)/H$ in Eq. (2.1). From a physical point of view, the term in parenthesis, which can be considered analogous to a peculiar velocity, reconciles the two approaches, based on the overdensity and radius, respectively.
From now on, we will neglect the contribution of shear and rotation in our analysis and set them to zero, effectively considering an exact spherical symmetry, as we want to focus on the effects of friction and avoid an interplay among them and friction. The effects of shear and rotation on structure formation have been studied before for both smooth and clustering dark energy models and for different implementations, as already discussed in Sect. 2. We will not delve into them again, but we refer the reader to [45, 46, 47, 48, 49, 50, 51, 52, 53] for details.
In Sect. 3 we detail how to calculate the time dependence of $\eta$ with respect to time. Here, we simply summarise our findings. For the dynamical friction model of [67] we obtain $\eta=\eta_{0}\,H$ with $\eta_{0}$ constant, and for dissipative pressure friction we find $\eta=(\eta_{1}+\eta_{2}|P/P_{0}|^{\Gamma})\,H_{0}$, with $\eta_{1}$, $\eta_{2}$, and $\Gamma$ all constant.
3 Theoretical models of friction
@@
3.1 Evolution of dynamical friction
Here, we closely follow the works by [67] and [71], to which we refer for more details.
In a scenario with hierarchical structure formation, structures of size $R$ form around the local maxima of the primordial density field smoothed over a scale of size $R$ [72, 73, 74] [^6] The peaks represent the fluctuations with positive density contrast in a Gaussian random field. Their properties can be calculated analytically following the theory of Gaussian random fields as outlined in the seminal work of [72]. We note, however, that many works showed the development of non-Gaussianity during the early evolution of cosmic structures [75, 76, 77, 78, 79]. The Gaussian approximation is thus a very useful starting point to achieve analytical results which allow progressing in the subject and a clear understanding of the underlying physics.. In this system, a test particle experiences a gravitational field which consists of a part associated with the smooth global mass distribution and a stochastic one which originates from the particle number fluctuations and gives rise to a frictional force $-\eta\mathbf{v}$ where $\eta$ is the dynamical friction coefficient and $\mathbf{v}$ the macroscopic velocity.
@@ 3.1 Evolution of dynamical friction
where we used the fact that $m_{\rm a}n_{\rm a}=\rho_{\rm m}=\bar{\rho}_{\rm m}\Delta$, with $\Delta$ the average overdensity of the perturbation. In principle, $\Delta$ would depend on the virialization recipe and on the cosmological model. However, here for simplicity, we will assume that it is constant, of the order of $\mathcal{O}(100)$.
This result shows that $\eta\propto H$ and that the value of $\eta_{0}$ depends on the total number of peaks $N$ within a protocluster, decreasing when $N$ increases. As discussed in [67], $N$ is considered to be approximately constant for clusters. For $N_{\rm tot}=10^{3}$ we find $\eta_{0}\approx 3\times 10^{-2}$, while for $N_{\rm tot}=10^{4}$ we have $\eta_{0}\approx 4\times 10^{-3}$. Regardless of the exact value of $\eta_{0}$, the main conclusion is that $\eta_{0}\ll 1$.
Note also that this result is not in contrast with previous work, such as [67], since an EdS model was assumed there. In fact, for such a model, neglecting radiation, we have $H\propto a^{-3/2}$. In fact, [80] showed that it is possible to associate a relaxation timescale with dynamical friction in galaxy clusters, and this is of the order of Hubble time, i.e., $1/\eta\propto H^{-1}$.
Finally, rather than expressing $\eta$ in terms of the total number of peaks $N$, it is worth using the mass of the perturbation $M$. Recalling that $M=m_{\rm a}N$, it is straightforward to obtain
@@
where we assume $m_{\rm a}=10^{9}\,h^{-1}\,M_{\odot}$. It is evident that $\eta_{0}$ is smaller for massive objects than for low-mass perturbations.
Eq. 13 comes from Eqs. D3–D5 in [71]. Those equations come from the theory of stochastic forces in a gravitational field, as shown by [81]. The terms $m_{a}$ and $n_{a}$ that determine the value of $\eta$ are related to the theory of Gaussian stochastic fields, valid for generic structures like galaxies, clusters, etc. In the text, we refer to the “protoclusters” just to estimate the values of these parameters. Again, in [71], the equations and values quoted are used to study clusters of galaxies.
Following [67], one can show that the peaks of the local density field with central height $\nu\geq\nu_{\rm c}$, with $\nu_{\rm c}$ a critical threshold, contribute to dynamical friction. The number of these objects can be calculated under the condition that the peak radius $r_{\rm pk}(\nu\geq\nu_{\rm c})$ is negligible with respect to the average peak separation $n_{\rm a}^{-1/3}(\nu\geq\nu_{\rm c})$.
The values required for the computation will necessarily depend on the particular cosmological model considered and on the filtering scale $R$. Here we do not pretend to derive exact values for the quantities above for every individual model we investigate in this work. Our aim is merely to estimate the order of magnitude of the relevant quantities, in particular the dynamical friction coefficient $\eta$ and to understand how this parameter affects the evolution of linear and nonlinear structures in $\Lambda$CDM and dark energy models.
Dynamical friction can be reinterpreted as a consequence of the fact that general relativity considers that a pressure term is not negligible within the dynamics dictated by the Friedmann equations. Hence, even if in Newtonian dynamics it is also possible to plug into the constraint equation for $H$ a pressure term, in fulfilment of energy conservation, the dynamical friction scenario ensures that friction is parametrised through the pressure itself, namely guaranteeing that the effects due to pressure are not subdominant than those inferred from the density alone. This generalises the previous scenario, where only density is involved, and proposes a way to better frame friction in time invoking that, from the general relativity recipe, the pressure effects are not negligible with respect to the density ones.
3.2 Dissipative pressure friction
Standard fluids can be grouped into two main classes: Newtonian and non-Newtonian. Clearly, the latter is found when the friction coefficient, $\eta$, depends on the variation of the shear velocity with respect to the radial coordinate. Our approach is simplified by virtue of the cosmological principle, i.e., the velocity field does not depend on radial coordinate. However, in this scenario, our friction coefficient is only a function of the Hubble rate. In other words, we do not emphasise the time dependence of it, assuming the simplest stationary case. It is clearly possible to generalise this case, including the effects due to the dynamics of the universe. Following the standard treatment, one can assume that the friction coefficient is a function of temperature and/or of thermodynamic quantities such as pressure, entropy, enthalpy and so forth. In the simplest case, we can assume what follows
@@ 3.2 Dissipative pressure friction
where $\eta_{1},\eta_{2}$ and $\Gamma$ are free real constants, with $P$ the total pressure of the universe and $P_{0}$ the total pressure today. The overall factor $H_{0}$ is just a normalisation factor that makes the two coefficients $\eta_{1}$ and $\eta_{2}$ dimensionless. From the second Friedmann equation,
@@
At this stage, we distinguish two cases, as reported in the following.
@@
    reducing to our previous case, discussed above, with $\Gamma=\tfrac{1}{2}$, $\eta_{1}=0$ and $\eta_{2}=\eta_{0}$. This certifies that our previous approach turns out to be a suitable approximation that works well in a phase in which the effects of pressure are subdominant over those associated with the density only over the dynamics of the universe, although it does not account for the whole stages of its evolution.
@@
Following the standard treatment for gases, one can assume that the friction coefficient may become a function of pressure, resembling the temperature or entropy dependence of quantities as for standard gases. Even though this hypothesis appears valid for real gases, we here enable it for cold dark matter, under the hypothesis that its contribution remains small and does not modify significantly the approximation of collisionless dark matter. For the sake of completeness, in fact, dark matter is not fully collisionless, see e.g. [82], and we may therefore expect a non-vanishing contribution provided by parameterizing the friction coefficient through the pressure as follows
@@
recasting in terms of the jerk parameter, defined as $j\equiv-H^{-1}\dot{q}+2q^{2}+q$ [24], with $q\equiv-1-\frac{\dot{H}}{H^{2}}$, the deceleration parameter and where we assumed that friction increases at large $z$ [^7] Quite remarkably, as a thermodynamic fact, we ensure friction to increase with redshift in order to guarantee that at early times friction might be more relevant in analogy to standard thermodynamic systems, where collisions appear more relevant as pressure and temperature increase. Nevertheless, since the universe temperature tends to larger values, one presumes that friction increases as well, as a consequence of the stronger interactions among constituents. Clearly, the interaction itself appears subdominant with respect to the underlying fluids, mainly dominated by radiation for very large redshifts. Accordingly, even if radiation tends to dominate over the other fluids, the friction effects are not zero, evolving up to current time, where their contribution is no longer negligible, as matter and radiation magnitude dilute, becoming smaller than at early times, around $z\simeq 0$.
The jerk parameter has been widely investigated in the literature [22, 29, 33, 34], adding more information about the way the Hubble parameter would evolve w.r.t. cosmic time. In particular, in order to guarantee the Universe to pass through a matter- to dark energy- dominated phases, one requires the jerk parameter to be always positive at late times. This may be seen evidently from the results obtained within the contexts of dark energy model-independent reconstructions, see e.g. [83, 84], where this has been established. As a matter of remark, this occurrence allows us to conclude $\mathrm{d}H/\mathrm{d}t>0$ but not that dark energy today turns into a cosmological constant [85]. Indeed, this may be argued instead if the jerk derivative would be well constrained, although this is currently not possible with current data.
Thus, since $\dot{H}>0$ and $j\geq 0$ [86, 87], we have
@@
where $\eta_{1}$ is arbitrarily imposed to vanish as it appears unessential in determining $\eta$ and the values ±$\pm 1$ selected for normalizing $\eta_{2}$ and $\Gamma$. These priors are not directly arguable from observations, but rather they match our current expectations over the cosmographic series inferred at late times, see e.g. [88, 89].
Having the above, case $\Gamma=0$ corresponds to a constant friction term, while case $\eta_{2}=0$ implies a frictionless model. Consequently, we argue that
@@
    the second case described above implies a model where friction acts repulsively, i.e., the sign is compatible with the dark energy equation of state. We will not consider this case here, as it does not represent the ordinary behavior of friction on matter.
@@
4 Evolution of linear perturbations